# Open Packing in Graphs: Bounds and Complexity

### v1

@@ 3.2 Subclasses of $(P_{4}\cup rK_{1})$-free Graphs
Let $G$ be a $(P_{4}\cup rK_{1})$-free graph. If $G$ does not contain any vertex subset that induces a $(P_{4}\cup(r-1)K_{1})$, then $G$ is $(P_{4}\cup(r-1)K_{1})$-free and so by induction assumption $\rho^{o}(G)\leq 2(r-1)+1<2r+1$. So, assume that $G$ has a vertex subset $D=\{x_{1},x_{2},x_{3},x_{4},y_{1},y_{2},\ldots,y_{r-1}\}$ that induces a $(P_{4}\cup(r-1)K_{1})$ in $G$ with $x_{i}x_{i+1}\in E(G)$ for $i=1,2,3$. Since $G$ is $(P_{4}\cup rK_{1})$-free, every vertex in $V(G)\setminus D$ is adjacent to some vertex in $D$, i.e., $D$ is a dominating set in $G$. So, $V(G)=N[x_{1}]\cup N[x_{2}]\cup N[x_{3}]\cup N[x_{4}]\cup N[y_{1}]\cup N[y_{2}]\cup\ldots\cup N[y_{r-1}]$. Hence, for any open packing $S$ in $G$, $S=\bigcup\limits_{u\in D}(S\cap N[u])$. Note that $|S\cap N[u]|\leq|S\cap N(u)|+|S\cap\{u\}|\leq 2$. Also, since $x_{1},x_{2},x_{3},x_{4}\in N(x_{2})\cup N(x_{3})$, $\bigcup\limits_{i=1}^{4}N[x_{i}]=\bigcup\limits_{i=1}^{4}N(x_{i})$. So, $|S\cap(\bigcup\limits_{i=1}^{4}N[x_{i}])|=|S\cap(\bigcup\limits_{i=1}^{4}N(x_{i}))|=|\bigcup\limits_{i=1}^{4}(S\cap N(x_{i}))|\leq 4$. Therefore, $|S|=|S\cap V(G)|=|S\cap(N[x_{1}]\cup N[x_{2}]\cup N[x_{3}]\cup N[x_{4}]\cup N[y_{1}]\cup N[y_{2}]\cup\ldots\cup N[y_{r-1}])|=|\bigcup\limits_{i=1}^{4}(S\cap N(x_{i}))|+|\bigcup\limits_{j=1}^{r-1}(S\cap N[y_{j}])|\leq 4+2(r-1)=2r+2$. Thus, to prove the claim, it is enough to rule out the possibility of $|S|=2r+2$. On the contrary, assume that there exists an open packing $S$ of size $22r+$ 2 in $G$. Then, by the above arguments, we can conclude that $|S\cap N[y_{j}]|=2$ for $j=1,2\ldots,r-1$ and $|S\cap(\bigcup\limits_{i=1}^{4}N(x_{i}))|=4$. Since $|S\cap N(y_{i})|\leq 1$ and $|S\cap N[y_{i}]|=2$, $y_{i}\in S$ and $S\cap N(y_{i})\neq\emptyset$. Let $S\cap N[y_{j}]=\{w_{j},y_{j}\}$ for $j=1,2,\ldots,r-1$, and let $S\cap(\bigcup\limits_{i=1}^{4}N(x_{i}))=\{z_{1},z_{2},z_{3},z_{4}\}$ with $x_{i}z_{i}\in E(G)$ (then, since $|S\cap N(x_{i})|\leq 1$, we have $S\cap N(x_{i})=\{z_{i}\}$ for every $i=1,2,3,4$). So, $S=\left(\bigcup\limits_{i=1}^{4}\{z_{i}\}\right)\cup\left(\bigcup\limits_{j=1}^{r-1}\{y_{j}\}\right)\cup\left(\bigcup\limits_{j=1}^{r-1}\{w_{j}\}\right)$. This together with the fact that $|S|=2r+2$ imply that every vertex in the $(2r+2)$-tuple $(z_{1},z_{2},z_{3},z_{4},y_{1},y_{2},\ldots,y_{r-1},w_{1},w_{2},\ldots,w_{r-1})$ is distinct. This implies that for $s\in\{1,2,3,4\}$, $z_{s}u\in E(G)$ for some $u\in D$ if and only if $u=x_{s}$ because of the fact that for every $u\in D$, (i) $S\cap N(u)=\{z_{i}\}$ if $u=x_{i}$ for $i\in\{1,2,3,4\}$ and (ii) $S\cap N(u)=\{w_{j}\}$ if $u=y_{j}$ for $j\in\{1,2,\ldots,r-1\}$. We complete our claim in two cases based on whether $z_{1}z_{4}\in E(G)$ or not.
@@
Therefore, $\rho^{o}(G)\leq 2r+1$. ∎
@@
The bound given in Lemma 8 is tight. For example, Let $G_{r}$ be a graph obtained by subdividing exactly $r$ edges of $K_{1,r+1}$ by four vertex paths. Graph $G_{3}$ is shown in Fig. 4. The vertex and edge set of $G_{r}$ can be defined as $V(G_{r})=(\cup_{i=1}^{r}\{x_{i},y_{i},z_{i}\})\cup\{u,v\}$ and $E(G_{r})=(\cup_{i=1}^{r}\{x_{i}y_{i},y_{i}z_{i},z_{i}u\})\cup\{uv\}$. Then, $G_{r}$ is ($P_{4}\cup rK_{1}$)-free (proof of this statement is given as Claim 11 in Appendix B) and $S_{r}=(\cup_{i=1}^{r}\{x_{i},y_{i}\})\cup\{v\}$ is an open packing in $G$ of size $12r+$ 1. Also, note that the graph obtained by replacing the vertex $u$ in $G_{r}$ with a complete graph will remain as a $(P_{4}\cup rK_{1})$-free graph with open packing number $12r+$ 1. Further, $\gamma_{t}(G_{r})=\rho^{o}(G)=2r+1$ with $D=\{u,z_{1},z_{2},\ldots,z_{r},y_{1},y_{2},\ldots,y_{r}\}$ as a total dominating set in $G_{r}$.
@@
In this section, we prove that for $r\geq 1$, $\gamma_{t}(G)\leq 2r+1$ and $\rho^{o}(G)\leq 2r$ for every connected ($P_{3}\cup rK_{1}$)-free graph $G$. Further, we show that the bound on (i) the open packing number in this graph class is tight and (ii) the total domination number is tight for the class of $(P_{3}\cup K_{1})$-free graphs.
@@
Every connected $P_{3}$-free graph is a complete graph and hence $\rho^{o}(G)\leq\gamma_{t}(G)=2$.
@@
Firstly, we prove the case $r=1$. Let $G$ be a connected ($P_{3}\cup K_{1}$)-free graph. If $G$ is $P_{3}$-free, then $\gamma_{t}(G)=2$. So, assume that $G$ contains a $P_{3}$. Let $D=\{x_{1},x_{2},x_{3}\}$ be a subset of $V(G)$ that induces a $P_{3}$ in $G$ with $x_{1}x_{2},x_{2}x_{3}\in E(G)$. Since $G$ is ($P_{3}\cup K_{1}$)-free, every vertex in $V(G)\setminus D$ is adjacent to some vertex in $D$. Further, since the vertices in $\{x_{1},x_{2},x_{3}\}$ are adjacent to a vertex in $\{x_{1},x_{2}\}\subseteq D$, every vertex in $V(G)=(V(G)\setminus D)\cup D$ is adjacent to some vertex in $D$. Hence, $D$ is a total dominating set in $G$ and $\gamma_{t}(G)\leq|D|=3=2r+1$. Next, we prove the case $r\geq 2$ by induction on $r$.
@@
Proof of Lemma 10 is similar to the proof of Lemma 8. Yet, the complete proof is given in Appendix B. The corollary below follows from Lemmas 9 and 10, and the fact that for $r\geq 3$, every $rK_{1}$-free graph is a ($P_{3}\cup(r-2)K_{1}$)-free graph as well.
@@
Figure 5: A connected $7K_{1}$-free graph $H_{7}$ with an open packing (as well as total dominating set) $S_{7}=\{x_{1},x_{2},x_{3},$ $x_{4},x_{5},y_{1},y_{2},y_{3},y_{4},y_{5}\}$ of size $2(7-2)=10$. Note that $H_{7}$ is also a ($K_{2}\cup 6K_{1}$)-free graph.
@@
Let $G$ be a connected ($K_{2}\cup K_{1}$)-free graph. Then $G$ is also a connected $P_{4}$-free graph. So, $\rho^{o}(G)\leq\gamma_{t}(G)=2(=r+1)$.
@@
Then, $\max\{r+1,2(r-1)\}=2(r-1)$. We prove that $\rho^{o}(G)\leq 2(r-1)$. On the contrary, assume that there exists a connected ($K_{2}\cup rK_{1}$)-free graph $G$ with $\rho^{o}(G)>2(r-1)$. We know that every open packing of size at least $12(r-1)+1=2r-$ 1 contains an open packing of size $12r-$ 1. So to contradict the assumption, it is enough to eliminate the possibility of the existence of an open packing of size $12r-$ 1 in $G$. Let $S$ be an open packing of size $12r-$ 1 in $G$. Then, since $G[S]$ is $\{P_{3},K_{3}\}$-free, every component of $G[S]$ contains at most two vertices and so the number of components in $G[S]$ is at least $\lceil\frac{2r-1}{2}\rceil>r-1$. Also, since $12r-$ 1 is odd, there exists at least one isolated vertex in $G[S]$ by the fact that every component of $G[S]$ contains at most two vertices. Let $u$ be the isolated vertex in $G[S]$. Since $G$ is connected, there exists a vertex $v\in V(G)$ such that $uv\in E(G)$. Note that $G[S\setminus\{u\}]$ contains at least $(r-1)$ components. Let $G_{1},G_{2},\ldots,G_{p}$ be the components of $G[S\setminus\{u\}]$ for some $p\geq r-1$. We now contradict to our assumption in two cases based on the value of $p$.
Case 1: $p\geq r$.
@@
Let $v_{i}\in V(G_{i})$ for $i=1,2,\ldots,r$. Since each $v_{i}$ is from a distinct component of $G[S\setminus\{u\}]$, $v_{i}\neq v_{j}$ for $i,j\in\{1,2,\ldots,r\}$ and $i\neq j$. Also, since $v_{i}\in S\setminus\{u\}$, $v_{i}\neq u$ and $v_{i}u\notin E(G[S])\subseteq E(G)$ by the fact that $u$ is an isolated vertex in $G[S]$. This implies that $v_{i}\neq v$ for $i=1,2\ldots,r$ since $uv\in E(G)$. Further $v_{i}v\notin E(G)$ since $v\in N(u)$ and $N(u)\cap N(v_{i})$ is empty because $S$ is an open packing in $G$ with $u,v_{i}\in S$ for every $i=1,2,\ldots,r$. This implies that $\{u,v\}\cup\{v_{1},v_{2},\ldots,v_{r}\}$ induces a ($K_{2}\cup rK_{1}$) in $G$, a contradiction.
Case 2: $p=r-1$.
@@
Then, $G_{1},G_{2},\ldots,G_{r-1}$ are the components in $G[S\setminus\{u\}]$. Recall that $|V(G_{i})|\leq 2$ for every $11\leq i\leq r-$ 1. But, since $22r-2=|S\setminus\{u\}|=|\bigcup\limits_{i=1}^{r-1}V(G_{i})|\leq 2r-$ 2, every $G_{i}$ consists of exactly two vertices. Let $V(G_{i})=\{u_{i},v_{i}\}$ for $11\leq i\leq r-$ 1. Since $G$ is connected, there exists a shortest path joining $u_{1}$ and $u$. Let $x_{1}x_{2}\ldots x_{s}$ with $u=x_{1}$ and $u_{1}=x_{s}$ be a shortest path joining $u_{1}$ and $u$ in $G$ (see Fig 6). Since $S$ is an open packing in $G$ with $u,v_{1}\in S$, $N(u)\cap N(v_{1})$ is empty. So, $uu_{1}\notin E(G)$ since $u_{1}v_{1}\in E(G)$. Similarly, $N(u)\cap N(u_{1})$ is empty. This implies that $s\geq 4$. So, $u\notin\{x_{s-2},x_{s-1}\}$. We further divide this case into two cases based on $x_{s-1}=v_{1}$ or not.
Case 2.1: $x_{s-1}\neq v_{1}$.
@@
Then, since $x_{s-1}x_{s}=x_{s-1}u_{1}\in E(G)$, i.e., $x_{s-1}\in N(u_{1})$ and $S$ is an open packing with $u_{1},v_{i}\in S$, $x_{s-1}\notin N(v_{i})$ (i.e., $x_{s-1}v_{i}\notin E(G)$) for every $i=1,2\ldots,r-1$. For similar reasons, $x_{s-1}u,x_{s-1}u_{2}\notin E(G)$. The existence of $u_{2}$ is guaranteed by the fact that $r-1\geq 2$. Also, note that $uv_{i},u_{2}v_{\ell},v_{i}v_{j}\notin E(G[S])\subseteq E(G)$ for every $i,j,\ell\in\{1,2,\ldots,r-1\}$ and $i\neq j$, $\ell\neq 2$ (see Fig. 6). Thus, $\{u_{2},v_{2}\}\cup\{x_{s-1}\}\cup\{v_{1}\}\cup\{v_{3}\}\cup\{v_{4}\}\cup\ldots\cup\{v_{r-1}\}\cup\{u\}$ is a ($K_{2}\cup rK_{1}$) in $G$, a contradiction.
Case 2.2: $x_{s-1}=v_{1}$.
@@
Then, similar to the above case, it can be proved that $x_{s-2}u,x_{s-2}u_{i},x_{s-2}v_{2}\notin E(G)$ for every $i=1,2\ldots,r-1$. Also, $uu_{i},v_{2}u_{\ell},u_{i}u_{j}\notin E(G)$ for every $i,j,\ell\in\{1,2,\ldots,r-1\}$ and $i\neq j$, $\ell\neq 2$. So, $\{u_{2},v_{2}\}\cup\{x_{s-2}\}\cup\{u_{1}\}\cup\{u_{3}\}\cup\{u_{4}\}\cup\ldots\cup\{u_{r-1}\}\cup\{u\}$ is a ($K_{2}\cup rK_{1}$) in $G$, a contradiction. ∎
@@
The bound given in Lemma 11 is tight. For $r\geq 3$, let $H_{r+1}$ be the graph as defined in Remark 5. Then, $H_{r+1}$ is a ($K_{2}\cup rK_{1}$)-free graph and the set $S_{r+1}$ in Remark 5 is an open packing of size $2(r-1)$.
@@
Given a graph class $\mathcal{G}$, Total Dominating Set is in P on the class $\mathcal{G}$ if $\mathcal{G}$ is (a) ($P_{4}\cup rK_{1}$)-free graphs, (b) $(P_{3}\cup rK_{1})$-free graphs, (c) $(K_{2}\cup rK_{1})$-free graphs or (d) $rK_{1}$-free graphs.
The above theorem follows from Lemmas 3, 6, 7 and 9, and Corollaries 1 and 2. The graph classes $(P_{3}\cup rK_{1})$-free graphs, $(K_{2}\cup rK_{1})$-free graphs and $rK_{1}$-free graphs are explicitly stated in Theorem 7 for the sake of clarity.
@@
Given a graph class $\mathcal{G}$, if $\mathcal{G}$ is a (a) $(P_{4}\cup rK_{1})$-free graphs ($r\geq 0$), (b) $(P_{3}\cup rK_{1})$-free graphs ($r\geq 0$), (c) $(K_{2}\cup rK_{1})$-free graphs ($r\geq 1$) or (d) $rK_{1}$-free graphs ($r\geq 3$), then for every fixed $r$ and for every graph $G\in\mathcal{G}$, (i) there are polynomially many open packings in $G$, (ii) all open packings in $G$ can be found in polynomial time, and hence $\rho^{o}(G)$ can be found in polynomial time.
The above theorem follows from Lemmas 5, 6, 8, 10, 11, Corollary 2, Remarks 4, 6 and the fact that for a graph $G$ with components $G_{1},G_{2},\ldots,G_{k}$, $\rho^{o}(G)=\sum_{i=1}^{k}\rho^{o}(G_{i})$. Next, we prove Theorem 2.
@@
For $p\geq 4$, let $H$ be a graph on $p$ vertices. Then, Open Packing is polynomial time solvable on the class of $H$-free graphs if and only if $H\in\{pK_{1},(K_{2}\cup(p-2)K_{1}),(P_{3}\cup(p-3)K_{1}),(P_{4}\cup(p-4)K_{1})\}$ unless $NP=P$.

### v2

@@ 3.2 Subclasses of $(P_{4}\cup rK_{1})$-free Graphs
Let $G$ be a $(P_{4}\cup rK_{1})$-free graph. If $G$ does not contain any vertex subset that induces a $(P_{4}\cup(r-1)K_{1})$, then $G$ is $(P_{4}\cup(r-1)K_{1})$-free, and so by induction assumption, $\rho^{o}(G)\leq 2(r-1)+1<2r+1$. So, assume that $G$ has a vertex subset $D=\{x_{1},x_{2},x_{3},x_{4},y_{1},y_{2},\ldots,y_{r-1}\}$ that induces a $(P_{4}\cup(r-1)K_{1})$ in $G$ with $x_{i}x_{i+1}\in E(G)$ for $i=1,2,3$. Since $G$ is $(P_{4}\cup rK_{1})$-free, every vertex in $V(G)\setminus D$ is adjacent to some vertex in $D$, i.e., $D$ is a dominating set in $G$. So, $V(G)=N[x_{1}]\cup N[x_{2}]\cup N[x_{3}]\cup N[x_{4}]\cup N[y_{1}]\cup N[y_{2}]\cup\ldots\cup N[y_{r-1}]$. Hence, for any open packing $S$ in $G$, $S=\bigcup\limits_{u\in D}(S\cap N[u])$. Note that $|S\cap N[u]|\leq|S\cap N(u)|+|S\cap\{u\}|\leq 2$. Also, since $x_{1},x_{2},x_{3},x_{4}\in N(x_{2})\cup N(x_{3})$, $\bigcup\limits_{i=1}^{4}N[x_{i}]=\bigcup\limits_{i=1}^{4}N(x_{i})$. So, $|S\cap(\bigcup\limits_{i=1}^{4}N[x_{i}])|=|S\cap(\bigcup\limits_{i=1}^{4}N(x_{i}))|=|\bigcup\limits_{i=1}^{4}(S\cap N(x_{i}))|\leq 4$. Therefore, $|S|=|S\cap V(G)|=|S\cap(N[x_{1}]\cup N[x_{2}]\cup N[x_{3}]\cup N[x_{4}]\cup N[y_{1}]\cup N[y_{2}]\cup\ldots\cup N[y_{r-1}])|=|\bigcup\limits_{i=1}^{4}(S\cap N(x_{i}))|+|\bigcup\limits_{j=1}^{r-1}(S\cap N[y_{j}])|\leq 4+2(r-1)=2r+2$. Thus, to prove the claim, it is enough to rule out the possibility of $|S|=2r+2$. On the contrary, assume that there exists an open packing $S$ of size $22r+$ 2 in $G$. Then, by the above arguments, we can conclude that $|S\cap N[y_{j}]|=2$ for $j=1,2,\ldots,r-1$ and $|S\cap(\bigcup\limits_{i=1}^{4}N(x_{i}))|=4$. Since $|S\cap N(y_{i})|\leq 1$ and $|S\cap N[y_{i}]|=2$, $y_{i}\in S$ and $S\cap N(y_{i})\neq\emptyset$. Let $S\cap N[y_{j}]=\{w_{j},y_{j}\}$ for $j=1,2,\ldots,r-1$, and let $S\cap(\bigcup\limits_{i=1}^{4}N(x_{i}))=\{z_{1},z_{2},z_{3},z_{4}\}$ with $x_{i}z_{i}\in E(G)$ (then, since $|S\cap N(x_{i})|\leq 1$, we have $S\cap N(x_{i})=\{z_{i}\}$ for every $i=1,2,3,4$). So, $S=\left(\bigcup\limits_{i=1}^{4}\{z_{i}\}\right)\cup\left(\bigcup\limits_{j=1}^{r-1}\{y_{j}\}\right)\cup\left(\bigcup\limits_{j=1}^{r-1}\{w_{j}\}\right)$. This, together with the fact that $|S|=2r+2$, imply that every vertex in the $(2r+2)$-tuple $(z_{1},z_{2},z_{3},z_{4},y_{1},y_{2},\ldots,y_{r-1},w_{1},w_{2},\ldots,w_{r-1})$ is distinct. This implies that for $s\in\{1,2,3,4\}$, $z_{s}u\in E(G)$ for some $u\in D$ if and only if $u=x_{s}$ because for every $u\in D$, (i) $S\cap N(u)=\{z_{i}\}$ if $u=x_{i}$ for $i\in\{1,2,3,4\}$ and (ii) $S\cap N(u)=\{w_{j}\}$ if $u=y_{j}$ for $j\in\{1,2,\ldots,r-1\}$. We complete our claim in two cases based on whether $z_{1}z_{4}\in E(G)$ or not.
@@
As we have arrived at a contradiction in both cases, $\rho^{o}(G)\leq 2r+1$. ∎
@@
The bound given in Lemma 8 is tight. For example, let $G_{r}$ be a graph obtained by subdividing exactly $r$ edges of $K_{1,r+1}$ by four vertex paths. Graph $G_{3}$ is shown in Fig. 4. The vertex and edge set of $G_{r}$ can be defined as $V(G_{r})=(\cup_{i=1}^{r}\{x_{i},y_{i},z_{i}\})\cup\{u,v\}$ and $E(G_{r})=(\cup_{i=1}^{r}\{x_{i}y_{i},y_{i}z_{i},z_{i}u\})\cup\{uv\}$. Then, $G_{r}$ is ($P_{4}\cup rK_{1}$)-free (proof of this statement is given as Claim 11 in Appendix B), and $S_{r}=(\cup_{i=1}^{r}\{x_{i},y_{i}\})\cup\{v\}$ is an open packing in $G$ of size $12r+$ 1. Also, note that the graph obtained by replacing the vertex $u$ in $G_{r}$ with a complete graph will remain a $(P_{4}\cup rK_{1})$-free graph with open packing number $12r+$ 1. Further, $\gamma_{t}(G_{r})=\rho^{o}(G)=2r+1$ with $D=\{u,z_{1},z_{2},\ldots,z_{r},y_{1},y_{2},\ldots,y_{r}\}$ as a total dominating set in $G_{r}$.
@@
In this section, we prove that for $r\geq 1$, $\gamma_{t}(G)\leq 2r+1$, and $\rho^{o}(G)\leq 2r$ for every connected ($P_{3}\cup rK_{1}$)-free graph $G$. Further, we show that the bound on (i) the open packing number in this graph class is tight and (ii) the total domination number is tight for the class of $(P_{3}\cup K_{1})$-free graphs.
@@
Every connected $P_{3}$-free graph is a complete graph, and hence $\rho^{o}(G)\leq\gamma_{t}(G)=2$.
@@
Firstly, we prove the case $r=1$. Let $G$ be a connected ($P_{3}\cup K_{1}$)-free graph. If $G$ is $P_{3}$-free, then $\gamma_{t}(G)=2$. So, assume that $G$ contains a $P_{3}$. Let $D=\{x_{1},x_{2},x_{3}\}$ be a subset of $V(G)$ that induces a $P_{3}$ in $G$ with $x_{1}x_{2},x_{2}x_{3}\in E(G)$. Since $G$ is ($P_{3}\cup K_{1}$)-free, every vertex in $V(G)\setminus D$ is adjacent to some vertex in $D$. Further, since the vertices in $\{x_{1},x_{2},x_{3}\}$ are adjacent to a vertex in $\{x_{1},x_{2}\}\subseteq D$, every vertex in $V(G)=(V(G)\setminus D)\cup D$ is adjacent to some vertex in $D$. Hence, $D$ is a total dominating set in $G$, and $\gamma_{t}(G)\leq|D|=3=2r+1$. Next, we prove the case $r\geq 2$ by induction on $r$.
@@
Proof of Lemma 10 is similar to the proof of Lemma 8. However, the complete proof is given in Appendix B. The corollary below follows from Lemmas 9 and 10, and the fact that for $r\geq 3$, every $rK_{1}$-free graph is a ($P_{3}\cup(r-2)K_{1}$)-free graph as well.
@@
Figure 5: A connected $7K_{1}$-free graph $H_{7}$ with an open packing (as well as a total dominating set) $S_{7}=\{x_{1},x_{2},x_{3},$ $x_{4},x_{5},y_{1},y_{2},y_{3},y_{4},y_{5}\}$ of size $2(7-2)=10$. Note that $H_{7}$ is also a ($K_{2}\cup 6K_{1}$)-free graph.
@@
Let $G$ be a connected ($K_{2}\cup K_{1}$)-free graph. Then, $G$ is also a connected $P_{4}$-free graph. So, $\rho^{o}(G)\leq\gamma_{t}(G)=2(=r+1)$.
@@
Then, $\max\{r+1,2(r-1)\}=2(r-1)$. We prove that $\rho^{o}(G)\leq 2(r-1)$. On the contrary, assume that there exists a connected ($K_{2}\cup rK_{1}$)-free graph $G$ with $\rho^{o}(G)>2(r-1)$. We know that every open packing of size at least $12(r-1)+1=2r-$ 1 contains an open packing of size $12r-$ 1. So, to contradict the assumption, it is enough to eliminate the possibility of an open packing of size $12r-$ 1 in $G$. Let $S$ be an open packing of size $12r-$ 1 in $G$. Then, since $G[S]$ is $\{P_{3},K_{3}\}$-free, every component of $G[S]$ contains at most two vertices, and so the number of components in $G[S]$ is at least $\lceil\frac{2r-1}{2}\rceil>r-1$. Also, since $12r-$ 1 is odd, there exists at least one isolated vertex in $G[S]$ due to the fact that every component of $G[S]$ contains at most two vertices. Let $u$ be the isolated vertex in $G[S]$. Since $G$ is connected, there exists a vertex $v\in V(G)$ such that $uv\in E(G)$. Note that $G[S\setminus\{u\}]$ contains at least $(r-1)$ components. Let $G_{1},G_{2},\ldots,G_{p}$ be the components of $G[S\setminus\{u\}]$ for some $p\geq r-1$. We now contradict our assumption in two cases based on the value of $p$.
Case 2.1: $p\geq r$.
@@
Let $v_{i}\in V(G_{i})$ for $i=1,2,\ldots,r$. Since each $v_{i}$ is from a distinct component of $G[S\setminus\{u\}]$, $v_{i}\neq v_{j}$ for $i,j\in\{1,2,\ldots,r\}$ and $i\neq j$. Also, since $v_{i}\in S\setminus\{u\}$, $v_{i}\neq u$ and $v_{i}u\notin E(G[S])\subseteq E(G)$ by the fact that $u$ is an isolated vertex in $G[S]$. This implies that $v_{i}\neq v$ for $i=1,2,\ldots,r$ since $uv\in E(G)$. Further $v_{i}v\notin E(G)$ since $v\in N(u)$ and $N(u)\cap N(v_{i})$ is empty because $S$ is an open packing in $G$ with $u,v_{i}\in S$ for every $i=1,2,\ldots,r$. This implies that $\{u,v\}\cup\{v_{1},v_{2},\ldots,v_{r}\}$ induces a ($K_{2}\cup rK_{1}$) in $G$, a contradiction.
Case 2.2: $p=r-1$.
@@
Then, $G_{1},G_{2},\ldots,G_{r-1}$ are the components in $G[S\setminus\{u\}]$. Recall that $|V(G_{i})|\leq 2$ for every $11\leq i\leq r-$ 1. But since $22r-2=|S\setminus\{u\}|=|\bigcup\limits_{i=1}^{r-1}V(G_{i})|\leq 2r-$ 2, every $G_{i}$ consists of exactly two vertices. Let $V(G_{i})=\{u_{i},v_{i}\}$ for $11\leq i\leq r-$ 1. Since $G$ is connected, there exists a shortest path joining $u_{1}$ and $u$. Let $x_{1}x_{2}\ldots x_{s}$ with $u=x_{1}$ and $u_{1}=x_{s}$ be a shortest path joining $u_{1}$ and $u$ in $G$ (see Fig 6). Since $S$ is an open packing in $G$ with $u,v_{1}\in S$, $N(u)\cap N(v_{1})$ is empty. So, $uu_{1}\notin E(G)$ since $u_{1}v_{1}\in E(G)$. Similarly, $N(u)\cap N(u_{1})$ is empty. This implies that $s\geq 4$. So, $u\notin\{x_{s-2},x_{s-1}\}$. We further divide this case into two cases based on whether $x_{s-1}=v_{1}$ or not.
Case 2.2.1: $x_{s-1}\neq v_{1}$.
@@
Then, since $x_{s-1}x_{s}=x_{s-1}u_{1}\in E(G)$, i.e., $x_{s-1}\in N(u_{1})$ and $S$ is an open packing with $u_{1},v_{i}\in S$, $x_{s-1}\notin N(v_{i})$ (i.e., $x_{s-1}v_{i}\notin E(G)$) for every $i=1,2,\ldots,r-1$. For similar reasons, $x_{s-1}u,x_{s-1}u_{2}\notin E(G)$. The existence of $u_{2}$ is guaranteed by the fact that $r-1\geq 2$. Also, note that $uv_{i},u_{2}v_{\ell},v_{i}v_{j}\notin E(G[S])\subseteq E(G)$ for every $i,j,\ell\in\{1,2,\ldots,r-1\}$ and $i\neq j$, $\ell\neq 2$ (see Fig. 6). Thus, $\{u_{2},v_{2}\}\cup\{x_{s-1}\}\cup\{v_{1}\}\cup\{v_{3}\}\cup\{v_{4}\}\cup\ldots\cup\{v_{r-1}\}\cup\{u\}$ is a ($K_{2}\cup rK_{1}$) in $G$, a contradiction.
Case 2.2.2: $x_{s-1}=v_{1}$.
@@
Then, similar to the above case, it can be proved that $x_{s-2}u,x_{s-2}u_{i},x_{s-2}v_{2}\notin E(G)$ for every $i=1,2,\ldots,r-1$. Also, $uu_{i},v_{2}u_{\ell},u_{i}u_{j}\notin E(G)$ for every $i,j,\ell\in\{1,2,\ldots,r-1\}$ and $i\neq j$, $\ell\neq 2$. So, $\{u_{2},v_{2}\}\cup\{x_{s-2}\}\cup\{u_{1}\}\cup\{u_{3}\}\cup\{u_{4}\}\cup\ldots\cup\{u_{r-1}\}\cup\{u\}$ is a ($K_{2}\cup rK_{1}$) in $G$, a contradiction. ∎
@@
The bound given in Lemma 11 is tight. For $r\geq 3$, let $H_{r+1}$ be the graph defined in Remark 5. Then, $H_{r+1}$ is a ($K_{2}\cup rK_{1}$)-free graph, and the set $S_{r+1}$ in Remark 5 is an open packing of size $2(r-1)$.
@@
Given a graph class $\mathcal{G}$, Total Dominating Set is in P on the class $\mathcal{G}$ if $\mathcal{G}$ is (a) ($P_{4}\cup rK_{1}$)-free graphs, (b) $(P_{3}\cup rK_{1})$-free graphs, (c) $(K_{2}\cup rK_{1})$-free graphs, or (d) $rK_{1}$-free graphs.
The above theorem follows from Lemmas 3, 6, 7, and 9, as well as Corollaries 1 and 2. The graph classes $(P_{3}\cup rK_{1})$-free graphs, $(K_{2}\cup rK_{1})$-free graphs, and $rK_{1}$-free graphs are explicitly stated in Theorem 7 for clarity.
@@
Given a graph class $\mathcal{G}$, if $\mathcal{G}$ is (a) $(P_{4}\cup rK_{1})$-free graphs ($r\geq 0$), (b) $(P_{3}\cup rK_{1})$-free graphs ($r\geq 0$), (c) $(K_{2}\cup rK_{1})$-free graphs ($r\geq 1$), or (d) $rK_{1}$-free graphs ($r\geq 3$), then for every fixed $r$ and for every graph $G\in\mathcal{G}$, (i) there are polynomially many open packings in $G$, (ii) all open packings in $G$ can be found in polynomial time, and hence $\rho^{o}(G)$ can be found in polynomial time.
The above theorem follows from Lemmas 5, 6, 8, 10, 11, Corollary 2, Remarks 4, 6, and the fact that for a graph $G$ with components $G_{1},G_{2},\ldots,G_{k}$, $\rho^{o}(G)=\sum_{i=1}^{k}\rho^{o}(G_{i})$. Next, we prove Theorem 2.